OpenStudy (anonymous):

hello... can you help me,to answer from the quetion below..and can you explains the basic theory.... question : A marble dropped ,which initially at rest,dropped on the oil box.This marble accelerated as function of velocity :A-Bv,with A and B are constant.find is velocity as time function.... thank's...

OpenStudy (anonymous):

thank's so much...

OpenStudy (anonymous):

\[a=A-Bv\] \[\int\limits_{0}^{v}dv/(A-Bv)=\int\limits_{0}^{t}dt\] Evaluating this gives the velocity as a function of time!! Easy!!